a. Suppose 1500 Joules of work is done on a system, and 300 Joules of heat are delivered to the system. What happens to the internal energy of that system?   

b. What is the efficiency of an engine that takes in heat at 200oC, and rejects heat to the environment at 25oC?   Big Hint: REMEMBER to work in Kelvin.

c. Bernoulli’s equation shows that pressure in a fluid will tend to drop as velocity increases.   Assuming a wind speed of 2 m/s causes a pressure drop of 0.700 Pa, what pressure drop is predicted by Bernoulli’s equation for a wind speed of 3 m/s?

d. How much heat is required to melt 75 g of ice?
Answer #1

Question a) first law :

ΔU = QW, where ΔU is the change in the internal energy, Q is the heat added to the system, and W is the work done by the system

Given W = -1500J and Q = 300 J

∆U = 300 -(-1500)

∆U = 1800 J

Question b)

Efficiency of engine = W/Qh

Th = 200°C = 473 K

Ti = 25°C = 298 K

Efficiency = Th - Ti/Th = 473 - 298/473 = 0.37

Efficieny (%) = 37%

Question C) p = (1/2) rho v^2.

0.7 = 1/2*rho*2^2

rho = 0.35

Since rho is constant

Pressure drop at 3m/s = 1/2*0.35*3^2 = 1.575 Pa

Question d) The heat of fusion of water is 333 J/g To melt 75 g

Heat required = 75* 333 = 24.975 KJ